OpenStudy (anonymous):

Solve for x : √8-2x=x+4

myininaya (myininaya):

So the radical is just over the 8?

myininaya (myininaya):

\[\sqrt{8}-2x=x+4? \] \[\sqrt{8-2x}=x+4 ?\]

OpenStudy (anonymous):

no its the second one

myininaya (myininaya):

Have you tried squaring both sides to get rid of the radical?

OpenStudy (anonymous):

I have no clue how to do that

myininaya (myininaya):

\[\text{ for example, we want to solve this one: } \sqrt{x}=9\] We square both sides \[(\sqrt{x})^2=(9)^2\] \[x=81 \]

myininaya (myininaya):

So you try squaring both sides for the one above

OpenStudy (anonymous):

uhh I have no clue , it sounds about right though

myininaya (myininaya):

Ok so when you have squared both sides you can draw it if you want what you will get after doing so

myininaya (myininaya):

So what did you get after squaring both sides?

myininaya (myininaya):

Here is another example Pretend we want to solve: \[\sqrt{x-2}=x+4\] We square both sides to get rid of the radical \[(\sqrt{x-2})^2=(x+4)^2\] \[(x-2)=(x+4)^2\] \[x-2=(x+4)^2\] Now remember the objective is to solve for x But we need to multiply that one squared junk out \[x-2=(x+4)(x+4)\] \[x-2=x^2+4x+4x+16\] \[x-2=x^2+8x+16\] Now this is a quadratic and I would put it in this form \[ax^2+bx+c=0 \text{ and use the quadratic formula } x=\frac{-b \pm \sqrt{b^2-4ac}}{2a} \text{ to solve }\]

OpenStudy (anonymous):

Ohhh , okay . so If I were to square both sides I would get (√8-2x)^2=(x-4)^2 ?

OpenStudy (anonymous):

And then 8-2x = x-4 ?

myininaya (myininaya):

well what happened to the square on the right hand side?

OpenStudy (anonymous):

ohh okay . so uhh 8-2x = (x-4) (x-4) right ?

myininaya (myininaya):

right! :)

myininaya (myininaya):

wait what happen to the plus

myininaya (myininaya):

in the orginal equation you had x+4 not x-4

OpenStudy (anonymous):

I just looked back at my probllem in my homework & I mistyped it , its actually x-4

myininaya (myininaya):

ok

OpenStudy (anonymous):

So thats it to the problem ? did I answer it ?

myininaya (myininaya):

No you have to solve for x

myininaya (myininaya):

Do you see in my example? The next step after what you have was to multiply (x-4)(x-4) out

myininaya (myininaya):

then I said to put in this form: \[ax^2+bx+c=0\] so you can use the quadratic formula

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

myininaya (myininaya):

That would be solving it for x because x is by itself

OpenStudy (anonymous):

Okay so my next step would be to 8-2=x^2-4x-4x-16 ?

myininaya (myininaya):

\[8-2x=x^2-4x-4x+16 \]

OpenStudy (anonymous):

oh snap , I was close ! lol okay so then it would be 8-2=x^2-8x+16 right ?

myininaya (myininaya):

you keep leaving that other x out but the other side is peachy! :)

myininaya (myininaya):

8-2x=x^2-8x+16

myininaya (myininaya):

Now subtract/add to both sides to put everything on one side

myininaya (myininaya):

You want everything on one side while 0 is on the other

OpenStudy (anonymous):

Ohh okay .. so what would i add or subtract first ;o

myininaya (myininaya):

No matters You just want everything on one side

OpenStudy (anonymous):

I got x^2+6x-8 , is that correct ?

myininaya (myininaya):

\[8-2x=x^2-8x+16 \] So I want everything on one side So I'm gonna add 2x on both sides And I'm gonna subtract 8 on both sides -8 +2x +2x -8 ----------------------- \[0=x^2-8x+2x+16-8\] \[0=x^2-6x+8\] Very close! :)

myininaya (myininaya):

Now you are almost done You could use the quadratic formula or try to factor (there is also completing the square but lets not-lol) Hint: factoring would work out nicely here actually

myininaya (myininaya):

Do you know how to factor?

OpenStudy (anonymous):

I think is it where you have to find to numbers that will multiply & add into the two numbers ?

myininaya (myininaya):

multiply to 8 and add to -6 :)

OpenStudy (anonymous):

4 & 2

myininaya (myininaya):

how about -4 and -2

myininaya (myininaya):

-4(-2)=8 -4+(-2)=-6

myininaya (myininaya):

so 0=(x-4)(x-2) :) Now set both factors =0

myininaya (myininaya):

But hey @breeskee03 you need to be careful about your solutions here

myininaya (myininaya):

because 8-2x has to be positive in order for there to be a solution

myininaya (myininaya):

So let me know what you get after solving 0=(x-4)(x-2) then I will go further with you okay?

OpenStudy (anonymous):

SO I plug it in now ?

myininaya (myininaya):

Set both factors =0 x-4=0 or x-2=0 Now solve both of these And then after you get those make sure when you plug into your original equation they both work

myininaya (myininaya):

You should be able to add 4 on both sides for x-4=0 And You should be able to add 2 on both sides for x-2=0

OpenStudy (anonymous):

okay (: Thank you so much !

myininaya (myininaya):

You got it?

myininaya (myininaya):

Thanks for working so hard to understand. I really like that about students. I'm proud of you for getting through this question. Go you! :)